OpenStudy (anonymous):

For the function below, provide the following: a. x-intercept(s) b. y-intercept(s) c. end behavior of the graph d. graph of this function. Please show all of your work f(x) = (x+5)^3

OpenStudy (anonymous):

x intercept is (0+5)^3=125 ---> (0,125)

OpenStudy (anonymous):

x intercept is (x+5)^3=0 (x+5)=cubed root 0=0 x+5=0 x=-5 So x intercept is (-5,0) The first i made I mistake it is the y intercept

OpenStudy (anonymous):

I am getting confused how you would graph this equation since it bounces at x=-5

OpenStudy (anonymous):

ok I graphed it

OpenStudy (anonymous):